# GLASS Flows: Transition Sampling for Alignment of Flow and Diffusion Models

### v1

@@ 6.1 Efficient Posterior Sampling and Value Function Estimation
Posterior sampling. We begin by benchmarking the efficiency of sampling transitions with GLASS Flows (our method) vs SDEs (DDPM sampling). As an example transition of particular importance, we use the posterior $p_{1|t}$ of the probability path (see eq. 9). We use DiT/SiT models from (Peebles & Xie, 2023; Ma et al., 2024), a competitive class-conditional flow matching model, trained on ImageNet256. Our experimental setup is as follows: We sample data points from the ImageNet model ($z\sim p_{\rm{data}}$), noise them ($x\sim p_{t}(\cdot|z)$), and then sample from the posterior via each respective method ($z^{\prime}\sim p_{1|t}(\cdot|x)$). For many simulation steps ($M=200$ in algorithm 1), both GLASS Flows and DDPM sampling give high quality samples from the posterior (see app. fig. 8). We then vary the number of simulation steps $M$ to values ranging from $M=2$ to $M=50$ and the time $t$. Note that the lower $M$ and the lower the time $t$, the “harder” the task gets as we have more discretization error and have added more noise to the reference image. Figures 9, 10 and 11 show that GLASS Flows return significantly higher quality samples for low $M$ or $t$. To quantify this, we measure the image quality via Frechet Inception Distance (FID) using 50k images for both reference and each method (for each combination of $t$ and $M$). GLASS Flows achieve significantly better FID than DDPM sampling for the same number of sampling steps (see fig. 2). Therefore, GLASS Flows represent a significant boost in efficiency when sampling from the posterior $p_{1|t}$.
@@ 6.3 Sequential Monte Carlo experiments
Table 1: Sequential Monte Carlo via Feynman-Kac steering (FKS). Every reward model defines a new experiment whose samples we evaluate on the same reward model and the GenEval benchmark. We set $N=8$ (number of particles). NFEs=$400$ for all rows except flow baseline ($50$ NFEs).
@@
Next, we apply GLASS Flows to inference-time reward alignment via Sequential Monte Carlo (SMC), in particular Feynman-Kac Steering (FKS) (Singhal et al., 2025; Skreta et al., 2025). We apply FKS on text-to-image generation using the FLUX model. We use a different pre-trained model than (Singhal et al., 2025) because FLUX is the current state-of-the-art model and our method requires a continuous-time model (the $t^{*}$ map does not fall into a discrete set of grid points). Except that, we use the same hyperparameters as in (Singhal et al., 2025). Between resampling steps, we sample the transitions with either DDPM sampling like previous works (Singhal et al., 2025) or GLASS Flows. We also compare against a Best-of-N baseline, i.e. where $N$ images are sampled and the one with highest reward selected. To ensure that we do not overfit to a single reward model, we run experiments for $4$ different reward models (CLIP (Hessel et al., 2021), Pick (Kirstain et al., 2023), HPSv2 (Wu et al., 2023b), ImageReward (Xu et al., 2023a)). Further, to avoid “reward hacking”, we evaluate results also on GenEval (Ghosh et al., 2023), to measure whether we can effectively optimize a reward without sacrificing GenEval performance. In table 1, we summarize our results and we plot examples in fig. 12. The first observation we make is that FKS with vanilla SDE sampling does not outperform a simple Best-of-N baseline as the Best-of-N is sampled with ODEs, i.e. the performance gain by using ODEs compared to SDEs weighs more than using SMC vs. Best-of-N. However, replacing the SDE transitions with GLASS Flows (our method), we remove this trade-off. In fact, GLASS Flows combined with FKS leads to significant improvements for all 4 rewards models without sacrificing performance on GenEval. We repeat the experiment on the PartiPrompts benchmark (Yu et al., 2022). Here, we optimize each reward model and evaluate on all other models. As shown in app. fig. 7, similarly FKS with GLASS Flows leads to significant improvements and also constitutes the best-performing method on PartiPrompts.
@@ 7 Conclusion
We introduce GLASS Flows, a novel way of sampling Markov transitions in flow and diffusion models using an “inner” flow matching model. This inner flow matching model can be retrieved from existing pre-trained models using sufficient statistics. Here, we applied GLASS Flows to inference-time reward alignment: Traditional sampling procedures for flow matching and diffusion models are poorly suited for inference-time reward alignment, either having deterministic trajectories like ODEs or requiring many steps to accurately simulate like SDEs. By combining the efficiency of ODEs with the stochasticity of SDEs, GLASS Flows substantially improve prior algorithms for inference-time reward alignment that relied on SDE sampling. Hence, GLASS Flows serve as a simple, drop-in solution for inference-time scaling of flow and diffusion models. In the future, we can apply GLASS Flows to other methods relying on SDE sampling, e.g. some reward fine-tuning (Domingo-Enrich et al., 2024) or image editing methods (Meng et al., 2021; Nie et al., 2023).
Acknowledgments
We thank Rafał Karczewski and Yutong He for helpful discussions. We thank all members of the Machine Learning and Inference Group at MIT (Tommi Jaakkola lab) for feedback on the draft, in particular Julia Balla, Abhi Gupta, Hannes Stärk, Felix Faltings, Cai Zhou, and Chenyu Wang. Further, we thank Benedikt Geiger, Alexander Sauer, and Rafał Karczewski for feedback on earlier versions of the draft.
@@ B.1 GLASS Flows and reward guidance
We explain how guidance aimed at sampling from the reward-tilted distribution as defined in Section 3 can be simply applied to GLASS Flows. Recall that to sample from the reward-tilted distribution $p^{r}(x)$ in our setting, the tilted vector field $u_{t}^{r}(x)$ can be written in terms of marginal vector field $u_{t}(x)$ and the value function $V_{t}(x)$
@@
$V_{t}(x)\approx\beta_{t}r(D_{t}(x))=:r_{t}(x)$
where $r_{t}$ as in eq. 10 and $\beta_{t}\geq 0$ is a hyperparameter (theoretically, $\beta_{t}=1$ would be ideal, it is common to tune this hyperparameter however). To derive a similar approximation for GLASS Flows, assume we have a pre-trained tilted denoiser $D_{t}^{r}(x)$ and consider applying GLASS Flows to the reward-tilted data distribution. Applying theorem 1 gives that the tilted GLASS velocity field can be written in terms of the tilted GLASS denoiser
$u^{r}_{s}(\bar{x}_{s}|x_{t},t)=w_{1}(s)\bar{x}_{s}+w_{2}(s)D_{\mu(s),\Sigma(s)}^{r}(x_{t},\bar{x}_{s})+w_{3}(s)x_{t},$ (45)
Further, proposition 2 gives that the tilted GLASS denoiser is given by
$D_{\mu,\Sigma}^{r}(\mathbf{x})=D_{t^{*}}^{r}\left(\alpha_{t^{*}}S(\mathbf{x})\right),$ (46)
where $D_{t}^{r}(x)$ is known by assumption. We can apply the same approximation to GLASS Flows by simply leveraging this approximate tilted denoiser as a replacement for the true $D_{t}^{r}(x)$, exactly as in approximate guidance for ODE’s or SDE’s. Specifically, we set
$u^{r}_{s}(\bar{x}_{s}|x_{t},t)=$ $u_{s}(\bar{x}_{s}|x_{t},t)+\beta_{t^{*}}\frac{\sigma_{t^{*}}^{2}}{\alpha_{t^{*}}}w_{2}(s)\nabla_{y}r(D_{t^{*}}(y))_{|y=\alpha_{t^{*}}S(\mathbf{x})}$ (47)
@@ B.2 $M=1$ GLASS Flows
As described in Section 4.3, GLASS Flows generate a data point using $K$ transitions and $M$ simulation steps per transition. For $K=1$, GLASS Flows are equal to standard flow matching integration of an ODE performed over $M$ simulation steps. In this section, we instead consider when $M=1$. Let ${\epsilon}\sim\mathcal{N}(0,I_{d})$, where $\bar{x}_{0}=\bar{\gamma}x_{t}+\bar{\sigma}_{0}{\epsilon}$. Then for one-step integration using the CondOT schedulers, we get

### v2

@@ 6.1 Efficient Posterior Sampling and Value Function Estimation
Posterior sampling. We begin by benchmarking the efficiency of sampling transitions with GLASS Flows (our method) vs SDEs (DDPM sampling). As an example transition of particular importance, we use the posterior $p_{1|t}$ of the probability path (see eq. 9). We use DiT/SiT models from (Peebles & Xie, 2023; Ma et al., 2024), a competitive class-conditional flow matching model, trained on ImageNet256. Our experimental setup is as follows: We sample data points from the ImageNet model ($z\sim p_{\rm{data}}$), noise them ($x\sim p_{t}(\cdot|z)$), and then sample from the posterior via each respective method ($z^{\prime}\sim p_{1|t}(\cdot|x)$). For many simulation steps ($M=200$ in algorithm 1), both GLASS Flows and DDPM sampling give high quality samples from the posterior (see app. fig. 10). We then vary the number of simulation steps $M$ to values ranging from $M=2$ to $M=50$ and the time $t$. Note that the lower $M$ and the lower the time $t$, the “harder” the task gets as we have more discretization error and have added more noise to the reference image. Figures 11, 12 and 13 show that GLASS Flows return significantly higher quality samples for low $M$ or $t$. To quantify this, we measure the image quality via Frechet Inception Distance (FID) using 50k images for both reference and each method (for each combination of $t$ and $M$). GLASS Flows achieve significantly better FID than DDPM sampling for the same number of sampling steps (see fig. 2). Therefore, GLASS Flows represent a significant boost in efficiency when sampling from the posterior $p_{1|t}$.
@@ 6.2 Novel Sampling Methods
In section D.3, we perform an ablation experiment over various correlation parameters $\rho$. Overall, we find that all values of $\rho$ are numerically stable and lead to ODE-level performance with only minor differences. The choice of a constant correlation schedule of $\rho=0.4$ led to the best results for the FLUX model and we use this in subsequent experiments. We note that the optimal choice of $\rho$ is dependent on the data and model and may well differ for other settings.
@@ 6.3 Sequential Monte Carlo experiments
Table 1: Sequential Monte Carlo via Feynman-Kac steering (FKS). Every reward model defines a new experiment whose samples we evaluate on the same reward model and the GenEval benchmark. We set $N=8$ (number of particles). NFEs=$400$ for all rows except flow baseline ($50$ NFEs). BoN: Best-of-N. FKS: Feynman-Kac Steering.
@@
Next, we apply GLASS Flows to inference-time reward alignment via Sequential Monte Carlo (SMC), in particular Feynman-Kac Steering (FKS) (Singhal et al., 2025; Skreta et al., 2025). We apply FKS on text-to-image generation using the FLUX model. We use a different pre-trained model than (Singhal et al., 2025) because FLUX is the current state-of-the-art model and our method requires a continuous-time model (the $t^{*}$ map does not fall into a discrete set of grid points). Except that, we use the same hyperparameters as in (Singhal et al., 2025). Between resampling steps, we sample the transitions with either DDPM sampling like previous works (Singhal et al., 2025) or GLASS Flows. We also compare against a Best-of-N baseline (BoN), i.e. where $N$ images are sampled and the one with highest reward selected. To ensure that we do not overfit to a single reward model, we run experiments for $4$ different reward models (CLIP (Hessel et al., 2021), Pick (Kirstain et al., 2023), HPSv2 (Wu et al., 2023b), ImageReward (Xu et al., 2023a)). Further, to avoid “reward hacking”, we evaluate results also on GenEval (Ghosh et al., 2023), to measure whether we can effectively optimize a reward without sacrificing GenEval performance. In table 1, we summarize our results and we plot examples in fig. 14. The first observation we make is that FKS with vanilla SDE sampling does not outperform a simple Best-of-N baseline as the Best-of-N is sampled with ODEs, i.e. the performance gain by using ODEs compared to SDEs weighs more than using SMC vs. Best-of-N. However, replacing the SDE transitions with GLASS Flows (our method), we remove this trade-off. In fact, GLASS Flows combined with FKS leads to significant improvements for all 4 rewards models without sacrificing performance on GenEval. We repeat the experiment on the PartiPrompts benchmark (Yu et al., 2022). Here, we optimize each reward model and evaluate on all other models. As shown in app. fig. 9, similarly FKS with GLASS Flows leads to significant improvements and also constitutes the best-performing method on PartiPrompts.
@@ 7 Conclusion
We introduce GLASS Flows, a novel way of sampling Markov transitions in flow and diffusion models using an “inner” flow matching model. This inner flow matching model can be retrieved from existing pre-trained models using sufficient statistics. Here, we applied GLASS Flows to inference-time reward alignment: Traditional sampling procedures for flow matching and diffusion models are poorly suited for inference-time reward alignment, either having deterministic trajectories like ODEs or requiring many steps to accurately simulate like SDEs. By combining the efficiency of ODEs with the stochasticity of SDEs, GLASS Flows substantially improve prior algorithms for inference-time reward alignment that relied on SDE sampling. Hence, GLASS Flows serve as a simple, drop-in solution for inference-time scaling of flow and diffusion models. In the future, one could explore applying GLASS Flows to other methods relying on SDE sampling, e.g. some reward fine-tuning (Liu et al., 2025; Xue et al., 2025; Li et al., 2025a; Domingo-Enrich et al., 2024) or image editing methods (Meng et al., 2021; Nie et al., 2023). Further, one could explore learning or dynamically adjusting the correlation parameter $\rho$ defining the GLASS transitions.
@@ B.1 GLASS Flows and reward guidance
We explain how (gradient) guidance aimed at sampling from the reward-tilted distribution as defined in Section 3 can be simply applied to GLASS Flows. We first recall the construction for “standard” FM and diffusion models and then show how to translate it to guidance for GLASS Flows.
Guidance for “standard” FM models.
Recall that to sample from the reward-tilted distribution $p^{r}(x)$ in our setting, the tilted vector field $u_{t}^{r}(x)$ can be written in terms of marginal vector field $u_{t}(x)$ and the value function $V_{t}(x)$
@@
$V_{t}(x)\approx\beta_{t}r(D_{t}(x))$ (45)
where $r_{t}$ as in eq. 10 and $\beta_{t}\geq 0$ is a hyperparameter (theoretically, $\beta_{t}=1$ would be ideal, it is common to tune this hyperparameter however). Therefore, the final approximated guidance vector is given by
$u_{t}^{r}(x)=u_{t}(x)+c_{t}\beta_{t}\nabla_{x}[r(D_{t}(x))]$
Guidance for GLASS Flows.
To derive guidance for GLASS Flows, we now translate the same principles to GLASS Flows. For this, let $D_{\mu,\Sigma}^{r}(\mathbf{x})$ be the denoiser for the reward-tilted distribution. Then we know that:
$D_{\mu(s),\Sigma(s)}^{r}(\mathbf{x})=D_{t^{*}}^{r}\left(\alpha_{t^{*}}S(\mathbf{x})\right)$ (46)
Further, using the same approximation in eq. 45 and inserting it into eq. 44, we get:
$D_{\mu(s),\Sigma(s)}^{r}(\mathbf{x})=$ $D_{t^{*}}^{r}\left(\alpha_{t^{*}}S(\mathbf{x})\right)$
$\approx$ $D_{t^{*}}\left(\alpha_{t^{*}}S(\mathbf{x})\right)+\beta_{t^{*}}\frac{\sigma_{t^{*}}^{2}}{\alpha_{t^{*}}}\nabla_{y}r(D_{t^{*}}(y))_{|y=\alpha_{t^{*}}S(\mathbf{x})}$
$=$ $D_{\mu(s),\Sigma(s)}(\mathbf{x})+\beta_{t^{*}}\frac{\sigma_{t^{*}}^{2}}{\alpha_{t^{*}}}\nabla_{y}r(D_{t^{*}}(y))_{|y=\alpha_{t^{*}}S(\mathbf{x})}$
Finally, we can insert this identity into the formula for the tilted GLASS velocity field (see theorem 1):
$u^{r}_{s}(\bar{x}_{s}|x_{t},t)$ (47)
$=$ $w_{1}(s)\bar{x}_{s}+w_{2}(s)D_{\mu(s),\Sigma(s)}^{r}(x_{t},\bar{x}_{s})+w_{3}(s)x_{t}$ (48)
$=$ $\underbrace{w_{1}(s)\bar{x}_{s}+w_{2}(s)D_{\mu(s),\Sigma(s)}(x_{t},\bar{x}_{s})+w_{3}(s)x_{t}}_{=u_{s}(\bar{x}_{s}|x_{t},t)}+\beta_{t^{*}}\frac{\sigma_{t^{*}}^{2}}{\alpha_{t^{*}}}w_{2}(s)\nabla_{y}r(D_{t^{*}}(y))_{|y=\alpha_{t^{*}}S(\mathbf{x})}$ (49)
$=$ $u_{s}(\bar{x}_{s}|x_{t},t)+\beta_{t^{*}}\frac{\sigma_{t^{*}}^{2}}{\alpha_{t^{*}}}w_{2}(s)\nabla_{y}r(D_{t^{*}}(y))_{|y=\alpha_{t^{*}}S(\mathbf{x})}$ (50)
where $u_{s}(\bar{x}_{s}|x_{t},t)$ is the GLASS velocity field for the (non-tilted) distribution $p_{\rm{data}}$. Theoretically, $\beta_{t}=1$ would be optimal for a perfect estimation of the value function. However, because of the approximation in eq. 45, we recommend tuning $\beta_{t}\geq 0$ as done already for previous guidance methods (Chung et al., 2022; He et al., 2023).
@@ B.2 $M=1$ GLASS Flows
As described in Section 4.2.2, GLASS Flows generate a data point using $K$ transitions and $M$ simulation steps per transition. For $K=1$, GLASS Flows are equal to standard flow matching integration of an ODE performed over $M$ simulation steps. In this section, we instead consider when $M=1$. Let ${\epsilon}\sim\mathcal{N}(0,I_{d})$, where $\bar{x}_{0}=\bar{\gamma}x_{t}+\bar{\sigma}_{0}{\epsilon}$. Then for one-step integration using the CondOT schedulers, we get
@@
B.3 Numerical stability
In the following, we show that is simple to ensure numerical stability in algorithm 1. Generally, we recommend performing all operations in algorithm 1 - except the neural network evaluation - in higher precision (float64). This has minimal overhead compared to large-scale neural network evaluations and minimize errors from reparameterization. We now discuss more specific steps.
$s=0$ edge case.
For $s=0$, it holds that
$\bar{X}_{0}=\bar{\gamma}X_{t}+\bar{\sigma}_{0}\epsilon$ (56)
One could think of this as a 2-step Markov chain, i.e. it holds $p(\bar{X}_{0}|X_{t},z)=p(\bar{X}_{0}|X_{t})$. It is a classical property of Markov chains that the posterior then also depends only on the state $X_{t}$:
$p(z|X_{t},\bar{X}_{0})=p(z|X_{t})$
One can prove this directly by applying Bayes’ rule twice and using the Markov property:
$p(z|X_{t},\bar{X}_{0})\propto$ $p(X_{t},\bar{X}_{0}|z)p(z)$ (57)
$=$ $p(\bar{X}_{0}|X_{t},z)p(X_{t}|z)p(z)$ (58)
$=$ $p(\bar{X}_{0}|X_{t})p(X_{t}|z)p(z)$ (59)
$\propto$ $p(X_{t}|z)p(z)$ (60)
$\propto$ $p(z|X_{t})$ (61)
As the first and last term both integrate to 1 (as they are probability distributions over $z$), they must be equal. As the posteriors are the same, also the denoisers are the same:
$D_{\mu(0),\Sigma(0)}(X_{t},\bar{X}_{0})=D_{t}(X_{t})$
In algorithm 1, we use this fact to ensure numerical stability at $s=0$.
Numerical stability of matrix inversion for $\Sigma(s)$ and weight coefficients.
The covariance matrix is given as:
$\Sigma(s)=\begin{bmatrix}\sigma_{t}^{2}&\sigma_{t}^{2}\bar{\gamma}\\ \sigma_{t}^{2}\bar{\gamma}&\bar{\sigma}_{s}^{2}+\bar{\gamma}^{2}\sigma_{t}^{2}\end{bmatrix}=\begin{bmatrix}\sigma_{t}^{2}&\sigma_{t^{\prime}}\sigma_{t}\rho\\ \sigma_{t^{\prime}}\sigma_{t}\rho&\bar{\sigma}_{s}^{2}+\rho^{2}\sigma_{t^{\prime}}^{2}\end{bmatrix}$ (62)
where we inserted $\bar{\gamma}=\rho\sigma_{t^{\prime}}/\sigma_{t}$. Then
$\det\Sigma(s)=(\bar{\sigma}_{s}^{2}+\rho^{2}\sigma_{t^{\prime}}^{2})\sigma_{t}^{2}-\sigma_{t^{\prime}}^{2}\sigma_{t}^{2}\rho^{2}=\sigma_{t}^{2}\bar{\sigma}_{s}^{2}$
Therefore, $\det\Sigma(s)>0$ and $\Sigma(s)$ is invertible whenever $\sigma_{t}^{2}>0$ and $\bar{\sigma}_{s}^{2}>0$. We now discuss when this might not be the case. First, $\sigma_{t}^{2}>0$ is equivalent to $t<1$ as $\sigma_{t}$ is strictly monotonically decreasing with $\sigma_{1}=0$ by assumption. Hence, $\sigma_{t}>0$ always holds in practice as we never take a transition starting at the final time $t=1$. However, it is important that the operation would not be well-defined in this case. Second, $\bar{\sigma}_{s}$ is also positive for $s<1$ by assumption and it fulfills at $s=1$ that $\bar{\sigma}^{2}_{1}=\sigma_{t^{\prime}}^{2}(1-\rho^{2})$. Therefore, for either $t^{\prime}=1$ or $\rho=\pm 1$, it would hold that $\bar{\sigma}_{1}=0$. Hence, for $t^{\prime}=1$ or $\rho=\pm 1$, the matrix $\Sigma(s)$ would not be invertible. However, in algorithm 1, we always simulate and take velocities for $s<1$. Therefore, everything is well-defined and we observed that the inversion of $\Sigma(s)$ did not constitute a numerical problem even for $\rho=\pm 1$. In fact, even for $\rho=\pm 1$, the samples we obtain are of high quality (see experiments in section D.3). Further, one can add a small value to the diagonal matrix to make it invertible: $\Sigma(s)\leftarrow\Sigma(s)+\epsilon I_{2}$ for $\epsilon>0$ to account for $s$ close to $1$. Similarly, the weight coefficients for the GLASS velocity field are given by:
for $\bar{\gamma}=\rho\sigma_{t^{\prime}}/\sigma_{t}$. We sample transitions for $t<t^{\prime}\leq 1$. Therefore, $t<1$ and also $\sigma_{t}>0$ and therefore $\bar{\gamma}$ is well-defined. Further, for $s=1$ and $\rho=\pm 1$ or $t^{\prime}=1$, it holds that $w_{1}(s)$ is not well-defined as $\bar{\sigma}_{s}=0$. However, as before, algorithm 1 only uses time steps $s\leq 1-1/M$ and therefore we did not encounter any numerical instabilities. As mentioned above, we recommend performing all operations in algorithm 1 - except the neural network evaluation - in higher precision (float64). This will have negligible overhead compared to neural network calls.
B.4 Other Diffusion Parameterizations
Other vector field parameterizations.
In algorithm 1, we assume that the pre-trained flow or diffusion model is given in the velocity parameterization $u_{t}(x)$ as used in flow matching. It is well-known that diffusion models can be equivalently parameterized via the score function $\nabla\log p_{t}(x)$ or the denoiser $D_{t}(x)$ or the noise predictor (also called $\epsilon$-predictor):
$\epsilon_{t}(x_{t})=\mathbb{E}[X_{0}|X_{t}=x_{t}]$
To use a model trained with a different parameterization for GLASS Flows, we simply reparameterize them into the denoiser parameterization $D_{t}(x)$. In algorithm 1, we have presented this for the reparameterization of $u_{t}$ into $D_{t}$ - for other models one might simply have to use the corresponding reparameterization. See e.g. (Lipman et al., 2024, Table 1) for reparameterization formulas.
Discrete-time parameterizations.
Our derivations assume that we have a model $u_{t}(x)$ trained in continuous time $0\leq t\leq 1$. In contrast, discrete-time diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are trained with a different time reparameterization. We discuss how to use GLASS Flows for these models. Let us assume that the discrete-time diffusion models is given in the shape of a denoiser model $\tilde{D}_{k}(x)$ with discrete time steps $k=1,\cdots,N$ and discrete-time schedulers $\tilde{\alpha}_{k},\tilde{\sigma}_{k}$. We can map these discrete-time $k=1,\cdots,N$ into a grid $\mathcal{G}=\{\tilde{t}_{j}\}_{j=1,\cdots,N}$ of continuous time points $0=\tilde{t}_{1}<\tilde{t}_{2}<\cdots<\tilde{t}_{N}=1$ via
$\tilde{t}_{k}=g^{-1}\left(\frac{\tilde{\sigma}_{k}^{2}}{\tilde{\alpha}_{k}^{2}}\right)$
where $g(t)=\sigma_{t}^{2}/\alpha_{t}^{2}$ as before for continuous-time schedulers $\alpha_{t},\sigma_{t}$. Further, define the denoiser model $D_{t}$ on the grid points as
$D_{\tilde{t}_{k}}(x)=\tilde{D}_{k}\left(\frac{\tilde{\alpha}_{k}}{\alpha_{t}}x\right)$
Note that $D_{t}(x)$ is a valid denoiser model for $t\in\mathcal{G}$ in the grid and schedulers $\alpha_{t},\sigma_{t}$ - same as before. However, there is one important difference: querying $D_{t}(x)$ for $t\notin\mathcal{G}$ would correspond to an invalid input to $\tilde{D}_{k}$ or, at least, and out-of-domain query of the neural network. Naturally, we want to avoid such out-of-domain queries. Specifically, during simulation of the GLASS Flow for a transition from $t$ to $t^{\prime}$, the denoiser model is queried at times $t^{*}(\mu(s),\Sigma(s))$ given by
$t^{*}(\mu(s),\Sigma(s))=$ $g^{-1}((\mu(s)\Sigma^{-1}(s)\mu(s))^{-1})$
$\mu(s)=$ $\begin{pmatrix}\alpha_{t}\\ \bar{\alpha}_{s}+\bar{\gamma}\alpha_{t}\end{pmatrix},\quad\Sigma(s)=\begin{pmatrix}\sigma_{t}^{2}&\sigma_{t}^{2}\bar{\gamma}\\ \sigma_{t}^{2}\bar{\gamma}&\bar{\sigma}_{s}^{2}+\bar{\gamma}^{2}\sigma_{t}^{2}\end{pmatrix}$
It holds that $t^{*}(\mu(0),\Sigma(0))=t$ (see section B.3) and therefore in particular we restrict transitions to only appear from grid points to grid points, i.e. $t,t^{\prime}\in\mathcal{G}$. To choose “inner” grid points $0=s_{0}<s_{1}<\cdots<s_{M}=1$, we can restrict ourselves to the set $\mathcal{T}=\{s\in[0,1]|t^{*}(\mu(s),\Sigma(s))\in\mathcal{G}\}$, i.e. choosing $s_{i}\in\mathcal{T}$. In general, there is no closed-form for $\mathcal{T}$ as we allow for general schedulers $\alpha_{t},\sigma_{t},\bar{\sigma}_{s},\bar{\alpha}_{s}$. A simple numerical approach to (approximately) obtain $\mathcal{T}$ is always valid. If $\alpha_{t},\sigma_{t},\bar{\sigma}_{s},\bar{\alpha}_{s}$ have simple analytical formulas, we might also obtain a closed form for $\mathcal{T}$. Therefore, GLASS Flows can be applied to discrete-time diffusion models in the same way and all of our results equally hold - with the only difference that the time points $s$ in algorithm 1 are constrained to $s\in\mathcal{T}$. Note that the above procedure is not an approximation but is exact: Simulating GLASS Flows with a reparameterized discrete-time diffusion model or a continuous-time diffusion model leads to identical results as long as time steps $s\in\mathcal{T}$ (assuming both models have no training error). Of course, there is still an error in the discretization of the simulation of the ODE - which would be identical for both, however. Therefore, GLASS Flows can also be applied to discrete-time diffusion models in the same way with a constrained set of valid inner grid points $s$ in algorithm 1.
B.5 Role of $\rho$
We briefly discuss the role of $\rho$ and how it determines the stochastic nature of the GLASS transition $p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t})$. Now that the GLASS transition can be written as:
$p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t})=\int\underbrace{p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t},z)}_{\text{GLASS (depends on }\rho)}\underbrace{p_{1|t}(z|x_{t})}_{\text{posterior (indep. of }\rho)}\mathrm{d}z$
Note note that for $\rho=1$, the conditional transition $p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t},z)$ becomes deterministic. However, $p_{1|t}(z|x_{t})$ is independent of $\rho$ and is not deterministic but rather a proper posterior distribution. Therefore, even for $\rho=1$, the GLASS transitions are stochastic (in particular, $\rho=1$ does not correspond to ODE sampling). Therefore, $\rho$ determines $p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t},z)$, in particular how correlated is that we add to a single data point but only partially determines the probabilistic/stochastic transition. We found that the time difference $t^{\prime}-t$ (determined by the number of transitions $K$) and the variance of the posterior $p_{1|t}$ are equally important factors determining the variance of the transition distribution $p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t})$.